\begin{thebibliography}{99}

\bibitem{AG19}
O. Ajbal and E. Godelle, \emph{Double centralizers in Artin--Tits groups},
Bull.\ Belg.\ Math.\ Soc.\ Simon Stevin \textbf{26} (2) (2019), 275--298.

\bibitem{Ar26}
E. Artin, \emph{Theorie der Z{\"o}pfe},
Abhandlungen aus dem Mathematischen Seminar der Hamburgischen Universit{\"a}t
\textbf{4} (1926), 47--72.

\bibitem{Ar47}
E. Artin, \emph{Theory of braids},
Annals of Math.\ (2) \textbf{48} (1947), 101--126.

\bibitem{BD99}
V. N. Bezverkhnii and I. V. Dobrynina,
\emph{On the unsolvability of the conjugacy problem for subgroups of the group
$R_5$ of pure braids},
Math.\ Notes \textbf{65} (1999), no.~1, 13--19.
Translated from Mat.\ Zametki \textbf{65} (1999), no.~1, 15--22.

\bibitem{Bo68}
N. Bourbaki,
\emph{\'{E}l\'{e}ments de math\'{e}matique. Fasc.\ XXXIV\@. Groupes et
alg\`ebres de Lie. Chapitre IV: Groupes de Coxeter et Syst\`emes de Tits.
Chapitre V: Groupes engendr\'es par des r\'eflexions. Chapitre VI:
Syst\`emes de racines},
Actualit\'{e}s Scientifiques et Industrielles, No.~1337,
Hermann, Paris, 1968.

\bibitem{DGM24}
F. Digne, E. Godelle and J. Michel,
\emph{Retraction to a parabolic subgroup and applications},
arXiv eprint 2407.07459, 2024.

\bibitem{GKLT13}
D. Garber, A. Kalka, E. Liberman and M. Teicher,
\emph{Double centralizers for parabolic subgroups of braid groups},
arXiv eprint 1310.0936, 2013.

\bibitem{Ga69}
F. A. Garside, \emph{The braid group and other groups},
Quart.\ J.\ Math.\ Oxford (2) \textbf{20} (1969), 235--254.

\bibitem{Go07}
E. Godelle, \emph{Parabolic subgroups of Garside groups},
J.\ Algebra \textbf{317} (2007), 1--16.

\bibitem{GW04}
J. Gonz{\'a}lez-Meneses and B. Wiest,
\emph{On the structure of the centralizer of a braid},
Ann.\ Sci.\ \'{E}cole Norm.\ Sup.\ (4) \textbf{37} (2004), no.~5, 729--757.

\bibitem{GrayNybergBrodda2025}
R. D. Gray and C.-F. Nyberg-Brodda,
\emph{Membership problems in braid groups and Artin groups},
J.\ London Math.\ Soc.\ \textbf{112} (2025), no.~4, article e70326.

\bibitem{Gu85}
G. Gurzo, \emph{Centralizers of finite sets of elements of the braid group
$B_{n + 1}$},
Mat.\ Zametki \textbf{37} (1985), 3--6.
English translation: Math.\ Notes \textbf{37} (1985), 3--4.

\bibitem{Ka07}
A. Kalka, \emph{Representations of braid groups and braid-based cryptography},
PhD thesis, Ruhr-Universit{\"a}t Bochum, 2007.

\bibitem{KLT10}
A. Kalka, E. Liberman and M. Teicher,
\emph{Solution to the subgroup conjugacy problem for Garside subgroups of
Garside groups},
Groups -- Complexity -- Cryptology \textbf{2}, no.~2 (2010), 157--174.

\bibitem{KTV14}
A. Kalka, B. Tsaban and G. Vinokur,
\emph{Complete simultaneous conjugacy invariants in Artin's braid groups},
arXiv eprint 1403.4622, 2014.

\bibitem{LL02}
S. J. Lee and E. Lee,
\emph{Potential weaknesses in the commutator key agreement protocol based on
braid groups},
Advances in Cryptology -- EUROCRYPT 2002, LNCS \textbf{2332},
Springer, 2002, 14--28.

\bibitem{lean}
A. Kalka, M. Teicher and B. Tsaban,
\emph{Lean verification code for this paper},
GitHub repository, commit \texttt{417828160c03b88eb7468f91bef025500c6aebcb},
\url{https://github.com/BoazTsaban/DCP-Lean-verification/tree/417828160c03b88eb7468f91bef025500c6aebcb}.

\bibitem{LyndonSchupp1977}
R. C. Lyndon and P. E. Schupp,
\emph{Combinatorial Group Theory},
Ergebnisse der Mathematik und ihrer Grenzgebiete, vol.~89,
Springer-Verlag, Berlin--New York, 1977.

\bibitem{Ma81}
T. A. Makanina,
\emph{The occurrence problem in the braid group $B_{n+1}$ when $n+1\ge 5$},
Mat.\ Zametki \textbf{29} (1981), no.~1, 31--33, 154.

\bibitem{Mi58}
K. A. Mikhailova,
\emph{The occurrence problem for direct products of groups},
Dokl.\ Akad.\ Nauk SSSR \textbf{119} (1958), 1103--1105. (Russian)

\bibitem{Mi66}
K. A. Mikhailova,
\emph{The occurrence problem for direct products of groups},
Mat.\ Sb.\ [Math.\ USSR-Sb.] \textbf{70}, no.~2 (1966), 241--251.

\bibitem{Miller1992}
C. F. Miller III,
\emph{Decision problems for groups---survey and reflections},
in: \emph{Algorithms and Classification in Combinatorial Group Theory},
Math.\ Sci.\ Res.\ Inst.\ Publ., vol.~23,
Springer, New York, 1992, pp.~1--59.

\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage{amssymb,amsmath} %, refcheck}
\usepackage[T1]{fontenc}
\usepackage[english]{babel}
\usepackage{cite}
\usepackage{tikz}
\usepackage{url}

\DeclareMathOperator{\acr}{cr} % algebraic crossing number

\newtheorem{proposition}{Proposition}[section]
\newtheorem{theorem}[proposition]{Theorem}
\newtheorem{maintheorem}[proposition]{Main Theorem}
\newtheorem{lemma}[proposition]{Lemma}
\newtheorem{mainlemma}[proposition]{Main Lemma}
\newtheorem{corollary}[proposition]{Corollary}

\theoremstyle{definition}
\newtheorem{definition}[proposition]{Definition}

\theoremstyle{remark}
\newtheorem{remark}[proposition]{Remark}

\newcommand{\inv}{^{-1}}

\begin{document}

\title[Double Coset Problem]{Double coset problem for parabolic subgroups of braid groups}
\author{Arkadius Kalka}
\author{Mina Teicher}
\author{Boaz Tsaban}
\address{Department of Mathematics, Bar-Ilan University, Ramat Gan, Israel}
\email{Arkadius.Kalka@rub.de}
\email{teicher@math.biu.ac.il}
\email{tsaban@math.biu.ac.il}

\begin{abstract}
We solve the double coset problem for all parabolic subgroups of braid groups.
The solution provides an effective reduction of this problem to the simultaneous conjugacy problem,
and resolves the ambiguity introduced by the center of the group
via the double centralizer theorem.
We also prove that the subgroup-restricted conjugacy problem is unsolvable in braid groups of at least 5 strands.
\end{abstract}

\subjclass[2020]{20F36; 20F10}

\keywords{braid groups, parabolic subgroups, double coset problem, simultaneous conjugacy problem, double centralizers}

\maketitle

\section{Double coset problem and other decision problems}\label{DCP}

Decision problems play an important role in combinatorial group theory~\cite{LyndonSchupp1977, Miller1992}.
Let $G$ be a group defined by generators and relations.
The \emph{word problem} in $G$ is the task of deciding whether a given group element,
provided as a product of group generators and their inverses, is the identity element.
The \emph{subgroup membership problem} is more general:
For a subgroup $H$ of $G$, decide whether a given group element $g$ lies in the subgroup $H$.
Another generalization of the word problem is the \emph{conjugacy problem}: Decide whether two given group elements are conjugate.
The \emph{simultaneous conjugacy problem} is to decide, given two lists of group elements $(g_1,\dotsc,g_k)$ and $(h_1,\dotsc,h_k)$, for some $k$, whether there is a group element $x$ such that $(x\inv g_1 x,\dotsc,x\inv g_k x)=(h_1,\dotsc,h_k)$.

The word problem and the conjugacy problem in braid groups have classical solutions, due to Artin~\cite{Ar26, Ar47} and Garside~\cite{Ga69}.
Lee and Lee~\cite{LL02} solved the simultaneous conjugacy problem in these groups,
and a later work of Kalka, Tsaban and Vinokur~\cite{KTV14} provides a more efficient solution.

Mikhailova~\cite{Mi58, Mi66} proved that the direct product $F_2\times F_2$, where $F_2$ is the free group with two generators,
has a finitely generated subgroup for which the subgroup membership problem is unsolvable.
Using this, and the existence of such direct products inside $B_n$ for $n\ge 5$,
Makanina~\cite{Ma81} proved that the subgroup membership problem for \emph{general} subgroups of the braid group $B_n$ is unsolvable for $n\ge 5$.

Gray and Nyberg-Brodda~\cite{GrayNybergBrodda2025} proved that the submonoid membership problem in $B_n$ is decidable
if and only if $n\le 3$.
In particular, the subgroup membership problem in $B_3$ is solvable.
The subgroup membership problem in $B_4$ remains open.

This paper focuses on the following decision problem.

\begin{definition}[Double Coset Problem]
	Let $A$ and $B$ be subgroups of a group $G$.
	Given an ordered pair $(g,h)\in G^2$, decide whether the element $h$ lies in the double coset $AgB$.
That is, whether there are elements $a\in A$ and $b\in B$ such that $h=agb$.
\end{definition}

The subgroup membership problem reduces to the double coset problem.
Indeed, given a subgroup $H$ of a group $G$, we have
$x\in H$ if and only if $x\in H\cdot 1\cdot \{1\}$.

Bourbaki~\cite{Bo68} solved the double coset problem for Coxeter groups.
Recently, Digne, Godelle and Michel~\cite{DGM24} obtained related results for the double coset problem in Artin groups beyond type $A$.
Kalka~\cite{Ka07} provided a reduction of a certain simultaneous double coset problem to the simultaneous conjugacy problem.
Gray and Nyberg-Brodda~\cite{GrayNybergBrodda2025} prove that the rational subset membership problem in the braid group $B_3$ is decidable.
Since the product of two rational subsets is rational,
double cosets $AgB$ of finitely generated subgroups are rational.
It follows that the double coset problem for finitely generated subgroups is solvable there.

\begin{corollary}
For $n\ge 5$, there are finitely generated subgroups $A$ and $B$ of the $n$-strand braid group for which the double coset problem is unsolvable.
\end{corollary}
\begin{proof}
The subgroup membership problem for a subgroup $H$
reduces to the double coset problem for the subgroups $H$ and $\{1\}$.
By Makanina's theorem~\cite{Ma81}, for $n\ge 5$ there exist finitely
generated subgroups $H$ of $B_n$ for which the membership problem is
unsolvable.
\end{proof}

It is thus natural to consider the double coset problem for important classes of subgroups of the braid group $B_n$.
Indeed, even the double coset problem for the naturally embedded braid subgroups $B_m\le B_n$ for $m<n$ was open prior to the present work.

\begin{definition}\label{def:parabolic}
	A subgroup $H$ of the braid group $B_n$ is \emph{standard parabolic} if
	it is generated by a subset of the set of standard generators $\{ \sigma_1, \dotsc, \sigma_{n-1}\}$.
	A subgroup $H$ of the braid group $B_n$ is \emph{parabolic} if
	it is conjugate to a standard parabolic subgroup.
\end{definition}

Our main result solves the double coset problem for all parabolic subgroups of $B_n$.

\begin{maintheorem}\label{thm:main}
Let $A$ and $B$ be parabolic subgroups of the braid group $B_n$.
The double coset problem for the pair $(A, B)$ in $B_n$ is solvable.
\end{maintheorem}

\begin{remark}
We assume that each parabolic subgroup is given by a standard parabolic subgroup
and a conjugating braid.
For mere decidability, this assumption is not necessary.
Indeed, given a finite set that generates a parabolic subgroup,
we may recover the needed data by a dovetailed (diagonal) enumeration:
enumerate subsets $I\subseteq\{1,\dotsc,n-1\}$, braids $\alpha\in B_n$,
and word certificates for the two inclusions between the given subgroup and
\[
\alpha\langle \sigma_i:i\in I\rangle\alpha\inv.
\]
Each proposed certificate is checked using the word problem in $B_n$.
Since the input subgroup is assumed to be parabolic, this search eventually
halts.
The question of computing the parabolic data efficiently is left open here.
\end{remark}

The proof proceeds by reduction of the double coset problem to an instance of the simultaneous conjugacy problem.
We use the following result~\cite[Theorem 3.16]{GKLT13}.
Ajbal and Godelle later established analogous double centralizer results for a broad class of Artin--Tits groups~\cite{AG19}.

\begin{theorem}[Garber, Kalka, Liberman and Teicher]\label{thm:double-centralizer}
	Let $H$ be a parabolic subgroup of the braid group $B_n$.
The double centralizer of the subgroup $H$ is given by
	\[ C_{B_n}(C_{B_n}(H))=\langle \Delta_n^2\rangle \cdot H. \]
\end{theorem}

For $n\ge 3$, the center $Z(B_n)$ is infinite cyclic, generated by $\Delta_n^2$,
where the braid $\Delta_n$ is the fundamental element.

\section{Decomposing parabolic modulo central elements}

Throughout this paper we assume $n\ge 3$;
the cases $n\le 2$ are immediate, since the group $B_1$ is trivial and the group $B_2$ is isomorphic to $\mathbb{Z}$.

\begin{definition}
Let $H=\langle \sigma_i : i\in I\rangle$ be a standard parabolic subgroup of the braid group $B_n$.
We denote by $\Pi_H$ the partition of $\{1,\dotsc,n\}$ into the connected components of the graph with vertex set $\{1,\dotsc,n\}$ and edge set $\{\{i,i+1\}:i\in I\}$.
\end{definition}

In other words, two strands lie in the same block of $\Pi_H$ if elements of $H$ may braid them with each other.
If a standard parabolic subgroup $H$ is proper, the partition $\Pi_H$ has at least two blocks.

\begin{proposition}\label{prp:decomposition}
Let $H$ be a proper standard parabolic subgroup of $B_n$, and $\alpha\in B_n$.
Let $G = \alpha H\alpha\inv $ be the associated parabolic subgroup.
Every element $x\in Z(B_n)\cdot G$ has a unique, effectively computable decomposition as a product of a central power and an element of $G$.
\end{proposition}
\begin{proof}
Let $x=\Delta_n^{2q}g$ with $q\in\mathbb Z$ and $g\in G$.
	Since the braid $\Delta_n^2$ is central, we have
\[
y := \alpha\inv x\alpha = \Delta_n^{2q}\alpha\inv g\alpha = \Delta_n^{2q}h,
\]
where $h := \alpha\inv g\alpha\in H$.

We are left with an instance of the problem
in a proper standard parabolic subgroup.
In this case,
the partition $\Pi_H$ has at least two blocks,
as illustrated in the following figure.
We fix strands, $r$ and $s$, in distinct components.

\begin{center}
	\begin{tikzpicture}[x=0.75cm,y=0.75cm,baseline=(current bounding box.center)]
		\foreach \i in {1,...,9} {
				\node[circle,fill=black,inner sep=1.5pt] (v\i) at (\i,0) {};
				\node[below=5pt] at (v\i) {\scriptsize $\i$};
			}
		\draw[thick] (v1)--(v2)--(v3);
		\draw[thick] (v5)--(v6);
		\draw[thick] (v8)--(v9);
		\draw[very thick] (v2) circle (4pt);
		\draw[very thick] (v8) circle (4pt);
		\node[above=9pt] at (v2) {\scriptsize $r$};
		\node[above=9pt] at (v8) {\scriptsize $s$};
	\end{tikzpicture}
\end{center}

	Erase from the braid $y$ all strands except the strands with left endpoints $r$ and $s$, and relabel the two remaining strands in their left-to-right order.

	Since the blocks of $\Pi_H$ are intervals and elements of $H$ braid only strands inside individual blocks, the two retained strands cannot cross in the image of the braid $h$.
Thus, the image of $h$ is the identity braid in the two-strand braid group $B_2$.

	The image of the braid $\Delta_n^{2q}$ is $\Delta_2^{2q}=\sigma_1^{2q}$.
	Thus, the erased braid is $\sigma_1^{2q}$, and the integer $q$ is effectively read off in $B_2\cong\mathbb Z$.
	We can thus compute the $G$-component as $g=\Delta_n^{-2q}x$.

This also shows that the power $2q$ is uniquely determined by the given element $x$, and thus the decomposition is unique.
\end{proof}

\section{Proof of the Main Theorem}
\label{Proof}

In this section we prove the Main Theorem.
The reduction produces an instance of the simultaneous conjugacy problem.
A solution of that instance is not, by itself, necessarily an element of the required double coset:
the double centralizer theorem leaves a possible central power.
The main point of the section is to show that this central power is effectively detected and, in the positive case, vanishes.

We first isolate the elementary combinatorial fact that supplies the detecting pair of strands.

\begin{lemma}\label{lem:detector-pair}
Let $A$ and $B$ be proper standard parabolic subgroups of the braid group $B_n$, and let $\pi$ be a permutation of $\{1,\dotsc,n\}$.
There are distinct strands $r$ and $s$ that lie in different blocks of $\Pi_A$ and whose images $\pi(r)$ and $\pi(s)$ lie in different blocks of $\Pi_B$.
\end{lemma}

\begin{proof}
Suppose, to the contrary, that every pair of strands from different blocks of $\Pi_A$ has its images in one block of $\Pi_B$.
Fix a block $P$ of $\Pi_A$ and a strand $s$ outside $P$.
For every $r\in P$, the strands $r$ and $s$ lie in different blocks of $\Pi_A$.
Thus $\pi(r)$ and $\pi(s)$ lie in the same block of $\Pi_B$.
Let this block be $Q$.
Then $\pi(P)\subseteq Q$.

Let $t\notin P$.
For any $r\in P$, the strands $r$ and $t$ lie in different blocks of $\Pi_A$.
Thus $\pi(r)$ and $\pi(t)$ lie in the same block of $\Pi_B$.
Since $\pi(r)\in Q$, we have $\pi(t)\in Q$.
It follows that $\pi(\{1,\dotsc,n\})\subseteq Q$.
This contradicts the fact that $B$ is proper, and hence that $\Pi_B$ has more than one block.
\end{proof}

For an element $u$ of the pure braid group $P_n$ and distinct strands $r$ and $s$, let $\acr_{r,s}(u)$ denote the algebraic crossing number of the two strands $r$ and $s$ in $u$.
Equivalently, after deleting all strands except $r$ and $s$, the resulting two-strand braid is $\sigma_1^{\acr_{r,s}(u)}$.
With this normalization, $\acr_{r,s}:P_n\to\mathbb Z$ is a homomorphism.
If $H$ is a standard parabolic subgroup, the strands $r$ and $s$ lie in different blocks of $\Pi_H$, and $u\in H\cap P_n$, then $\acr_{r,s}(u)=0$.
Also,
\[
\acr_{r,s}(\Delta_n^{2m})=2m
\]
for all $m\in\mathbb Z$ and all $r\ne s$.

We use the standard naturality of these homomorphisms under conjugation.
For every braid $g\in B_n$ there is a permutation $\rho_g$ of $\{1,\dotsc,n\}$ such that, for all $u\in P_n$,
\[
\acr_{r,s}(gug\inv)=\acr_{\rho_g(r),\rho_g(s)}(u).
\]

\begin{mainlemma}\label{lem:central-power}
Let $A$ and $B$ be proper parabolic subgroups of the $n$-strand braid group $B_n$.
If elements $a\in A$, $b\in B$, and $g\in B_n$ satisfy
\[
gbg\inv=\Delta_n^{2k}a,
\]
then $k=0$.
\end{mainlemma}

\begin{proof}
Choose braids $\alpha,\beta\in B_n$ such that
\[
A=\alpha A_0\alpha\inv,\qquad B=\beta B_0\beta\inv,
\]
where $A_0$ and $B_0$ are standard parabolic subgroups.
Write $a=\alpha a_0\alpha\inv$ and $b=\beta b_0\beta\inv$.
Since $\Delta_n^2$ is central, the equation $gbg\inv=\Delta_n^{2k}a$ is equivalent to
\[
(\alpha\inv g\beta)b_0(\alpha\inv g\beta)\inv=\Delta_n^{2k}a_0.
\]
Thus it is enough to prove the assertion when $A$ and $B$ are standard parabolic.

Assume, then, that $A$ and $B$ are standard parabolic.
Let $\nu\colon B_n\to S_n$ be the homomorphism induced by the action on strand endpoints, and let $p$ be the order of the permutation $\nu(a)$.
From $gbg\inv=\Delta_n^{2k}a$ we get
\[
\nu(g)\nu(b)\nu(g)\inv=\nu(a).
\]
Thus $\nu(b)$ also has order $p$.
Consequently, $a^p$ and $b^p$ are pure braids, and
\[
gb^p g\inv=\Delta_n^{2kp}a^p.
\]

Apply Lemma~\ref{lem:detector-pair} to the permutation $\rho_g$.
Choose strands $r$ and $s$ in different blocks of $\Pi_A$ such that $\rho_g(r)$ and $\rho_g(s)$ lie in different blocks of $\Pi_B$.
Since $b^p\in B\cap P_n$, naturality gives
\[
\acr_{r,s}(gb^p g\inv)
=\acr_{\rho_g(r),\rho_g(s)}(b^p)=0.
\]
On the other hand, since $a^p\in A\cap P_n$ and $r,s$ lie in different blocks of $\Pi_A$,
\[
\acr_{r,s}(\Delta_n^{2kp}a^p)
=
2kp+\acr_{r,s}(a^p)
=
2kp.
\]
Hence $2kp=0$.
Since $p\ge 1$, it follows that $k=0$.
\end{proof}

We now reduce the double coset problem to simultaneous conjugacy.

\begin{theorem}\label{thm:reduction}
Let $A$ and $B$ be parabolic subgroups of the braid group $B_n$.
The double coset problem for the pair $(A,B)$ effectively reduces to the simultaneous conjugacy problem in $B_n$.
\end{theorem}

\begin{proof}
Centralizers of parabolic subgroups of braid groups are finitely generated and effectively computable~\cite{Gu85,GKLT13}.
Let
\[
C_{B_n}(A)=\langle c_1,\dotsc,c_{k_A}\rangle,\qquad
C_{B_n}(B)=\langle d_1,\dotsc,d_{k_B}\rangle
\]
be such generating sets.

Let $g,g'\in B_n$ be an instance of the double coset problem for the pair $(A,B)$.
Thus we have to decide whether
\[
g'=agb
\]
for some $a\in A$ and $b\in B$.
If $A=B_n$ or $B=B_n$, then $AgB=B_n$, and the instance is positive.
We may therefore assume that both $A$ and $B$ are proper, as required by
Main Lemma~\ref{lem:central-power}.

Consider the following simultaneous conjugacy instance, with unknown $x$:
\begin{equation}\tag{$*$}\label{eq:sc-instance}
\left\{
\begin{aligned}
x\inv c_i x &= c_i
    && i=1,\dotsc,k_A, \\
x\inv(g' d_j {g'}\inv)x &= g d_j g\inv
    && j=1,\dotsc,k_B.
\end{aligned}
\right.
\end{equation}
This is an effectively constructed instance of the simultaneous conjugacy problem.
We use a solution $x$ when the instance is positive.

If~\eqref{eq:sc-instance} has no solution, then $g'\notin AgB$.
Indeed, if $g'=agb$ with $a\in A$ and $b\in B$, then $a$ satisfies all equations in~\eqref{eq:sc-instance}.
The first set holds because each $c_i$ centralizes $A$.
The second set holds because each $d_j$ centralizes $B$:
\[
g'd_j{g'}\inv
=agbd_jb\inv g\inv a\inv
=a(gd_jg\inv)a\inv.
\]

Assume now that~\eqref{eq:sc-instance} has a solution $x$.
The first set of equations says that $x$ centralizes every generator of $C_{B_n}(A)$.
Thus
\[
x\in C_{B_n}(C_{B_n}(A)).
\]
By the double centralizer theorem and Proposition~\ref{prp:decomposition}, compute the unique decomposition
\[
x=\Delta_n^{2k}a_0
\]
with $k\in\mathbb Z$ and $a_0\in A$.

Set
\[
y:=g\inv x\inv g'.
\]
The second set of equations gives, for every $j$,
\[
g\inv x\inv g'd_j{g'}\inv xg=d_j.
\]
Thus $y$ centralizes every generator of $C_{B_n}(B)$, and hence
\[
y\in C_{B_n}(C_{B_n}(B)).
\]
Again, by the double centralizer theorem and Proposition~\ref{prp:decomposition}, compute the unique decomposition
\[
y=\Delta_n^{2l}b_0
\]
with $l\in\mathbb Z$ and $b_0\in B$.
Since $g'=xgy$, we have
\[
g'=\Delta_n^{2(k+l)}a_0gb_0.
\]
If $k+l=0$, then $g'=a_0gb_0$, and the original double coset instance is positive.
It remains to show that, when the original instance is positive, this test cannot produce $k+l\ne 0$.

Suppose, then, that $g'=agb$ for some $a\in A$ and $b\in B$.
As noted above, $a$ is also a solution of~\eqref{eq:sc-instance}.
Since both $a$ and $x$ solve the first set of equations,
\[
a\inv x\in C_{B_n}(C_{B_n}(A))
=\langle \Delta_n^2\rangle\cdot A.
\]
Since both $a$ and $x$ solve the second set of equations,
\[
a\inv x\in gC_{B_n}(C_{B_n}(B))g\inv
=\langle \Delta_n^2\rangle\cdot gBg\inv.
\]
Thus, for some $\bar a\in A$, $\bar b\in B$, and $p,q\in\mathbb Z$,
\[
x=\Delta_n^{2p}a\bar a
=\Delta_n^{2q}ag\bar b g\inv.
\]
Comparing the two expressions gives
\[
g\bar b g\inv=\Delta_n^{2(p-q)}\bar a.
\]
By Main Lemma~\ref{lem:central-power}, $p=q$.

The decomposition of $x$ in $\langle\Delta_n^2\rangle\cdot A$ is unique, so $k=p$.
Moreover,
\[
y
=g\inv x\inv g'
=g\inv (g\bar b\inv g\inv a\inv \Delta_n^{-2q})agb
=\Delta_n^{-2q}\bar b\inv b.
\]
Thus, the central exponent of $y$ in $\langle\Delta_n^2\rangle\cdot B$ is $l=-q$.
Since $p=q$, we have
\[
k+l=p-q=0.
\]
It follows that the algorithm returns a positive answer precisely for the positive instances.
\end{proof}

\begin{proof}[Proof of Main Theorem~\ref{thm:main}]
For $n\le 2$, the braid group $B_n$ is either trivial or infinite cyclic, and the assertion is immediate.
Assume $n\ge 3$.
By Theorem~\ref{thm:reduction}, the double coset problem for parabolic subgroups of $B_n$ effectively reduces to the simultaneous conjugacy problem in $B_n$,
which is solvable~\cite{LL02}.
In a positive instance, the Lee--Lee algorithm~\cite{LL02} produces a
conjugator.
This completes the proof of Main Theorem~\ref{thm:main}.
\end{proof}

\section{The unsolvability of the subgroup-restricted conjugacy problem}

The double coset problem is related to the following problem.

\begin{definition}[Subgroup-restricted Conjugacy Problem]
Given a subgroup $H$ of a group $G$ and elements $a,b\in G$, decide whether there exists an element $c\in H$ such that
$b=c\inv ac$.
\end{definition}

This problem was previously named \emph{subgroup conjugacy problem}~\cite{KLT10, GKLT13}.
The present name distinguishes it from the \emph{conjugacy problem for subgroups},
where we are given two finitely generated subgroups $A,B\leq G$ and the question is whether
$g\inv Ag=B$ for some element $g\in G$.
The latter problem is unsolvable in the pure braid group $R_5$~\cite{BD99}.

Kalka, Liberman and Teicher~\cite{KLT10} solved this problem for parabolic subgroups of braid groups and, more generally,
for Garside subgroups of Garside groups in the sense of Godelle~\cite{Go07}.
They suggest that this problem may be unsolvable in $B_n$, for $n\ge 5$.
Later, Garber, Kalka, Liberman and Teicher~\cite{GKLT13} proved that for parabolic subgroups of $B_n$,
this problem is reducible to the simultaneous conjugacy problem.

\begin{theorem}\label{thm:subgroup-conjugacy-unsolvable}
For $n\ge 5$, there is a finitely generated subgroup $H\le B_n$
for which the subgroup-restricted conjugacy problem is unsolvable.
Consequently,
the subgroup-restricted conjugacy problem in $B_n$, for general finitely generated subgroups, is unsolvable.
\end{theorem}

\begin{proof}
Makanina's construction~\cite{Ma81} gives pure braids
\[
a=\sigma_3^2,\quad
b=\sigma_3\sigma_2^2\sigma_3,\quad
c=\sigma_4\sigma_3\sigma_2^2\sigma_3\sigma_4,\quad
d=\sigma_4\sigma_3\sigma_2\sigma_1^2\sigma_2\sigma_3\sigma_4
\]
such that $\langle a,b\rangle$ and $\langle c,d\rangle$ are free groups of rank two,
commute elementwise, and intersect trivially. Thus
\[
M=\langle a,b,c,d\rangle=\langle a,b\rangle\times \langle c,d\rangle
\]
is a subgroup of $P_n$ isomorphic to $F_2\times F_2$.
Here, for $n\ge 5$, the braids $a,b,c,d$ are regarded as elements of $B_n$ by adding trivial strands.
By Mikhailova's theorem~\cite{Mi58,Mi66}, there is a finitely generated subgroup
$L\le M$ for which the subgroup membership problem in $M$ is unsolvable.
We will reduce this membership problem to the subgroup-restricted conjugacy problem for the fixed subgroup $L\le B_n$.

Set
\[
\delta=\sigma_1\sigma_2\dotsc\sigma_{n-1}.
\]
As shown by Gonz{\'a}lez-Meneses Wiest~\cite{GW04},
the standard centralizer computation for periodic braids gives
\begin{equation}\label{eq:delta-centralizer}
C_{B_n}(\delta)=\langle \delta\rangle.
\end{equation}
We also have $\delta^n=\Delta_n^2$~\cite{Ga69}.

We claim that
\begin{equation}\label{eq:M-delta-trivial}
M\cap C_{B_n}(\delta)=1.
\end{equation}
Indeed, let $x\in M\cap C_{B_n}(\delta)$.
By~\eqref{eq:delta-centralizer}, $x=\delta^q$ for some $q\in\mathbb Z$.
Let $\nu\colon B_n\to S_n$ be the natural homomorphism.
Since $M\le P_n=\ker\nu$, we have $\nu(x)=1$.
But $\nu(\delta)$ has order $n$, and therefore $n$ divides $q$.
Thus $x=\Delta_n^{2r}$ for some $r\in\mathbb Z$.
The braid $\Delta_n^2$ is central in $B_n$.
Hence $x$ is central in $M$.
Since $M\cong F_2\times F_2$ has trivial center, $x=1$.
This proves~\eqref{eq:M-delta-trivial}.

Now let $w\in M$ be an instance of the membership problem for $L$.
Construct the following instance of the subgroup-restricted conjugacy problem for the subgroup $L$:
\[
\alpha=\delta,
\qquad
\beta=w\inv\delta w.
\]
We show that this instance is positive if and only if $w\in L$.
If $w\in L$, then $\beta=w\inv \alpha w$, and the element $w\in L$ is a valid conjugator.

Conversely, suppose that the instance is positive.
Then there is an element $\ell\in L$ such that
\[
w\inv\delta w=\ell\inv\delta \ell.
\]
Multiplying on the left by $w$ and on the right by $\ell\inv$, we get
\[
\delta w\ell\inv=w\ell\inv\delta.
\]
Thus $w\ell\inv\in C_{B_n}(\delta)$.
Since $w\in M$ and $\ell\in L\le M$, we also have $w\ell\inv\in M$.
By~\eqref{eq:M-delta-trivial}, $w\ell\inv=1$.
Therefore $w=\ell\in L$.

Thus, an algorithm for the subgroup-restricted conjugacy problem for $L\le B_n$ would decide the subgroup membership problem for $L\le M$, contradicting Mikhailova's theorem.
\end{proof}

\section{Discussion and open problems}

The proof above is modular, and its ingredients indicate what would be needed in order to extend the result beyond braid groups.
Let $H$ be a proper standard parabolic subgroup of a group $G$.
We need effective generators for the centralizers of the relevant parabolic subgroups, together with a double centralizer theorem of the form $C(C(H))=Z(G)\cdot H$.
We also need an effective way to decompose elements known to lie in the product $Z(G)\cdot H$ as products of central elements and elements of $H$.
In the braid-group case, the center is infinite cyclic, generated by $\Delta_n^2$.
An analogous argument in a group whose center is generated by a central element $z$ would also require a substitute for Main Lemma~\ref{lem:central-power}, namely that the equation $gbg\inv =z^k a$ forces $k=0$ whenever the elements $a$ and $b$ lie in proper standard parabolic subgroups $A$ and $B$, respectively.

This suggests extending Main Theorem~\ref{thm:main} to parabolic subgroups of other Artin--Tits groups, for example Artin groups of finite types $B$ and $D$.
Such an extension would require an algebraic replacement for the parts of the present proof that use braid-specific geometry, especially erasing strands and pairwise crossing numbers.
The same issue arises in generalizations of the Main Lemma.

Even within braid groups, the following problem remains open:
Are the subgroup membership problem and the double coset problem solvable in $B_4$?

\subsection*{Artificial intelligence and formal verification}
The authors used ChatGPT Pro for paper improvement, editorial assistance, and preliminary error-checking.
Codex and Lean 4 were used to formally verify the proofs,
modulo the cited black-box braid-theoretic and algorithmic theorems.
The Lean verification code is available on GitHub~\cite{lean}.

The Lean verification code for the finite computations used in this paper is
available at the GitHub repository cited in~\cite{lean}.  The cited commit is
the version checked for the present manuscript.

The verification formalizes the local arguments that do not require a Lean library for braid groups: the reduction from subgroup membership to the double coset problem, the detector-pair lemma, the correction step for the central ambiguity, and the integer arithmetic in Main Lemma~\ref{lem:central-power}.

\subsection*{Acknowledgments}
Much of this research was conducted when the first author was a postdoctoral researcher at Bar-Ilan University, hosted by the second and third authors.
He thanks them for their hospitality during this period.
The first author acknowledges financial support by the Minerva Foundation of Germany.
The second author acknowledges financial support by the Oswald Veblen Fund.

\end{document}